OpenStudy (anonymous):

Find the t required for a $1200 investment to double at interest rate r, compound continuously.

OpenStudy (amistre64):

what was wrong with the same question you posted in the general math section?

OpenStudy (anonymous):

tmi doubling time has nothing to do with the initial investment i invest $50 and @amistre64 invests $700,000 mine double to $100 and his doubles to $1,400,000 in the same amount of time it rains on the just and the unjust

OpenStudy (amistre64):

indeed. and we will need to recall the formula for continuous compounding interest rates

OpenStudy (anonymous):

just unsure how to solve for t

OpenStudy (amistre64):

you need to know the formula for continuous compounded interest. Then solving for t will be rather simple

OpenStudy (anonymous):

A=Pe^rt from what I know, but not that simple to me really...

OpenStudy (amistre64):

lets redefine A, we want it to be 2P right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

... fixing a few typos A=Pe^rt now lets start peeling away the layers to get to "t" 2P = Pe^rt ; divide each side by P 2 = e^rt ; ln undoes an e ln(2) = rt now it just a matter of removing the r ....

OpenStudy (anonymous):

well r is given as 0.062

OpenStudy (anonymous):

I'm sorry for not stating so from the start, slipped my mind

OpenStudy (amistre64):

can you solve: ln(2) = rt , for t?

OpenStudy (anonymous):

divide both sides by r?

OpenStudy (amistre64):

correct soo ln(2)/r = t use a calculator and plug in the value for r to determine t

OpenStudy (anonymous):

11.17

OpenStudy (anonymous):

and note that when @amistre64 writes 2P = Pe^rt ; divide each side by P 2 = e^rt ; it means that the initial amount is a red herring. doubling time is doubling time

OpenStudy (anonymous):

oh I see. Thank you so much